Carla Vista Monograms sells stadium blankets that have been monogrammed with high school and university emblems. The blanketssolve C te wwwwwwwww1.790 Carvalbar d 51000 w Ok.be www.buti Pep pers Text ook and Weda wur warmt. Andrea Wunscret. Car scret 1000 re
Answer #1

Selling price per unit = $49

variable cost per unit = 40% of selling price per unit

= 49 x 40%

= $19.6

Fixed cost = $120,000 per month

Proposed changes :

Variable cost per unit = 47% of selling price per unit

= 49 x 47%

= $23.03

Increase in foxed cost = $12,100 per month

Fixed cost (After increase) = 120,000 + 12,100

= $132,100 per month

Annual fixed cost = 132,100 x 12

= $1,585,200

Increase in selling price = 12%

= 49 x 12%

= $5.88

Selling price per unit (after increase) = 49 + 5.88

= $54.88

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 54.88 - 23.03

= $31.85

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 31.85/54.88

= 58.04%

Break even point in dollar = Fixed cost/Contribution margin ratio

= 1,585,200/58.04%

= $2,731,219.85

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.